# WorldGPT: Empowering LLM as Multimodal World Model

### v1

@@ 1. Introduction
With the capability to process any modality in any domain, WorldGPT can serve as a universal world simulator. Unlike previous generation models that only support simulating static scenes (e.g., Stable Diffusion (Rombach et al., 2022)) or image transformations based on naive editing instructions (e.g., InstructPix2Pix (Brooks et al., 2022)), WorldGPT is capable of synthesizing dynamic scenes that change according to complex interactions, offering more practical significance. Utilizing WorldGPT, we explore a novel learning paradigm for multimodal agents (e.g., MLLMs (Zhu et al., 2023a; Ye et al., 2023)), namely dream tuning, where agents acquire specialized knowledge from WorldGPT to enhance their performance on specific tasks by fine-tuning on synthetic multimodal instruction data. The generation process for this data is efficient: a powerful LLM (e.g., GPT-4) generates the textual components of instructions and drives WorldGPT to complete the multimodal part. We conducted experiments with four widely-used MLLM agents across three diverse tasks ((Visual Understanding, Embodied Planning, and Audio-Video Question Answering). Results reveal that agents trained on synthesized data exhibit competitive performance compared to those trained on authentic data, strongly supporting the reliability of WorldGPT as a world simulator.
@@ 2. Related Work
Multimodal Data Synthesis. While extensive studies focus on generating textual instructions for a broad range of tasks (Wang et al., 2023a; Peng et al., 2023; Yang et al., 2023c), the synthesis of multimodal instruction data is still confined to the following applications: 1) Given real images, generate textual instructions with LLM (Liu et al., 2023b; Lyu et al., 2023). For instance, LLaVA (Liu et al., 2023b) and Macaw-LLM (Lyu et al., 2023) pre-collect multimodal data from existing datasets and then prompt ChatGPT (Achiam et al., 2023) to generate textual instructions. 2) Given collected captions, generate images with generation models (Sharifzadeh et al., 2024; Fan et al., 2023). The synthesized image-caption pairs can then serve as reliable pretraining resources. 3) Given real images, apply an image editing model that takes specific types of instructions, such as color, style, and object change (Brooks et al., 2022; Li et al., 2023b). For example, VPG-C (Li et al., 2023b) synthesizes counterfactual images using Blended Diffusion (Avrahami et al., 2022) given original images and editing instructions.
@@ 3.1. MLLM as Foundation of World Model
For multimodal encoding, we first utilize LanguageBind to obtain a unified representation for the multimodal state. Then, similar with other MLLMs (Wu et al., 2023b; Han et al., 2023a), a linear projection layer will be applied to convert features from multimodal embedding into LLM embedding. To help LLM better utilize multimodal information, we introduce special tokens to specify the multimodal inputs. For example, a state represented by both video and audio will be transformed into the following sequence ”<VID> [video embedding] </VID> <AUD> [audio embedding] </AUD>” before sending to LLM.
For multimodal decoding, inspired by recent work (Zheng et al., 2023b), we include special tokens into LLM’s vocabulary as multimodal signals. When these special tokens have been decoded exhaustively, i.e., ”[<IMG1><IMG2>...<IMGn>]”, we decode the corresponding hidden states with a trainable transformer-based projection layer. In this paper, we investigate two types of target output space: original modality space and language-aligned feature space. For the first one, the output of transformer-based projection layer will be trained to aligned with corresponding conditional text representations of the diffusion models. This allows states predicted in abstract spaces to be accurately reconstructed in real-world modality spaces. Sometimes, we don’t need to reconstruct the features because they will be re-encoded into the abstract space for the next input step, such as for downstream multimodal agents. In this cases, the output of transformer-based projection layer will map the signal back to the unified multimodal encoding space (i.e. LanguageBind feature). Since most multimodal agents take the multimodal inputs using a similar language-aligned multimodal encoder (e.g., CLIP (Radford et al., 2021)), an applicable way would be applying a lightweight projector (e.g. linear layer) to connecting the output of WorldGPT to their input encoding space which avoids the cumbersome process of decoding to real-world process and encoded back again.
@@ 4.1. Details of the Cognitive Architecture
Knowledge retrieval system manages external knowledge, such as collections of state transitions, and supports retrieval actions through dense retrieval. All states and actions within these collections are encoded by a unified encoder, enabling efficient retrieval. When applied to a new domain, WorldGPT enhances its capabilities by retrieving the most similar experiences from the pre-collected knowledge base. For instance, consider a scenario where WorldGPT is employed as a simulator of a chemical laboratory, an environment replete with unique cases not encompassed in the training datasets, such as chemical reactions. If similar cases can be supplied by the knowledge retrieval system, WorldGPT could effectively predict outcomes by ”imitating” the retrieved samples.

### v2

@@ 1. Introduction
With the capability to process any modality in any domain, WorldGPT can serve as a universal world simulator. Unlike previous generation models that only support simulating static scenes (e.g., Stable Diffusion (Rombach et al., 2022)) or image transformations based on naive editing instructions (e.g., InstructPix2Pix (Brooks et al., 2022)), WorldGPT is capable of synthesizing dynamic scenes that change according to complex interactions, offering more practical significance. Utilizing WorldGPT, we explore a novel learning paradigm for multimodal agents (e.g., MLLMs (Zhu et al., 2023a; Ye et al., 2023)), namely dream tuning, where agents acquire specialized knowledge from WorldGPT to enhance their performance on specific tasks by fine-tuning on synthetic multimodal instruction data. The generation process for this data is efficient: a powerful LLM (e.g., GPT-4) generates the textual components of instructions and drives WorldGPT to complete the multimodal part. We conducted experiments with four widely-used MLLM agents across three diverse tasks (Visual Understanding, Embodied Planning, and Audio-Video Question Answering). Results reveal that agents trained on synthesized data exhibit competitive performance compared to those trained on authentic data, strongly supporting the reliability of WorldGPT as a world simulator.
@@ 2. Related Work
Multimodal Data Synthesis. While extensive studies focus on generating textual instructions for a broad range of tasks (Wang et al., 2023a; Peng et al., 2023; Yang et al., 2023c), the synthesis of multimodal instruction data (Yang et al., 2023c; Fei et al., 2024d) is still confined to the following applications: 1) Given real images, generate textual instructions with LLM (Liu et al., 2023b; Lyu et al., 2023). For instance, LLaVA (Liu et al., 2023b) and Macaw-LLM (Lyu et al., 2023) pre-collect multimodal data from existing datasets and then prompt ChatGPT (Achiam et al., 2023) to generate textual instructions. 2) Given collected captions, generate images with generation models (Sharifzadeh et al., 2024; Fan et al., 2023). The synthesized image-caption pairs can then serve as reliable pretraining resources. 3) Given real images, apply an image editing model that takes specific types of instructions, such as color, style, and object change (Brooks et al., 2022; Li et al., 2023b; Fei et al., 2024c). For example, VPG-C (Li et al., 2023b) synthesizes counterfactual images using Blended Diffusion (Avrahami et al., 2022) given original images and editing instructions.
@@ 3.1. MLLM as Foundation of World Model
For multimodal encoding, we first utilize LanguageBind to obtain a unified representation for the multimodal state. Then, similar with other MLLMs (Wu et al., 2023b; Han et al., 2023a; Fei et al., 2024d), a linear projection layer will be applied to convert features from multimodal embedding into LLM embedding. To help LLM better utilize multimodal information, we introduce special tokens to specify the multimodal inputs. For example, a state represented by both video and audio will be transformed into the following sequence ”<VID> [video embedding] </VID> <AUD> [audio embedding] </AUD>” before sending to LLM.
For multimodal decoding, inspired by recent work (Zheng et al., 2023b; Pan et al., 2024), we include special tokens into LLM’s vocabulary as multimodal signals. When these special tokens have been decoded exhaustively, i.e., ”[<IMG1><IMG2>...<IMGn>]”, we decode the corresponding hidden states with a trainable transformer-based projection layer. In this paper, we investigate two types of target output space: original modality space and language-aligned feature space. For the first one, the output of transformer-based projection layer will be trained to aligned with corresponding conditional text representations of the diffusion models. This allows states predicted in abstract spaces to be accurately reconstructed in real-world modality spaces (Fei et al., 2024b, a, d). Sometimes, we don’t need to reconstruct the features because they will be re-encoded into the abstract space for the next input step, such as for downstream multimodal agents. In this cases, the output of transformer-based projection layer will map the signal back to the unified multimodal encoding space (i.e. LanguageBind feature). Since most multimodal agents take the multimodal inputs using a similar language-aligned multimodal encoder (e.g., CLIP (Radford et al., 2021)), an applicable way would be applying a lightweight projector (e.g. linear layer) to connecting the output of WorldGPT to their input encoding space which avoids the cumbersome process of decoding to real-world process and encoded back again.
@@ 4.1. Details of the Cognitive Architecture
Knowledge retrieval system manages external knowledge, such as collections of state transitions, and supports retrieval actions through dense retrieval. All states and actions within these collections are encoded by a unified encoder, enabling efficient retrieval. When applied to a new domain (Pan et al., 2023), WorldGPT enhances its capabilities by retrieving the most similar experiences from the pre-collected knowledge base. For instance, consider a scenario where WorldGPT is employed as a simulator of a chemical laboratory, an environment replete with unique cases not encompassed in the training datasets, such as chemical reactions. If similar cases can be supplied by the knowledge retrieval system, WorldGPT could effectively predict outcomes by ”imitating” the retrieved samples.